OpenStudy (anonymous):

prove the identity: [(tanx)/(1-cosx) ] = [(cscx)(1+secx)]

OpenStudy (anonymous):

\[\left(\begin{matrix}tanx \\ 1-cosx\end{matrix}\right)\] = (cscx)(1+secx)

OpenStudy (cwrw238):

rewrite tanx as sin x / cos x and simplify should help

OpenStudy (cwrw238):

[(tanx)/(1-cosx) ] = sin x / ( cos x - cos^2 x) cscx (1 + sec x) = 1/sinx * ( 1 + 1/cosx) = 1/sinx * ( ( cos x + 1) / cos x) = ( cos x + 1) / sinx cos x - i think i'm on the wrong track here!!!

OpenStudy (cwrw238):

- there is probably a simpler way to do this

OpenStudy (anonymous):

\[\frac{tanx}{1-cosx}=cscx(1+secx)\] \[tanx=(1+secx-cosx-cosxsecx)cscx\] \[tanx=(secx-cosx)cscx\] \[tanx=(\frac{1}{cosx}-cosx)cscx\] \[tanx=(\frac{1-\cos^2x}{cosx})cscx\] \[tanx=\frac{\sin^2x cscx}{cosx}\] \[tanx=tanxsinxcscx\] \[tanx=tanx\]

OpenStudy (cwrw238):

good job

OpenStudy (anonymous):

ty